# Convergence Rates for Stochastic Proximal and Projection Estimators

### v2

@@ 4.1. Stochastic Approximation of Projections onto Smooth Sets
Let $C\subset\mathbb{R}^{n}$ be a closed convex set and let $p\in\partial C$. We say that $\partial C$ admits a local $C^{2,1}$ chart at $p$ with constants $(\rho,L,M)$ if, after a rigid motion sending $p$ to 0 and the outward unit normal at $p$ to $e_{n}$, there exists $\rho>0$ and a concave function $h\in C^{2,1}(\mathbb{B}^{n-1}_{\rho})$ such that $h(0)=0$, $\nabla h(0)=0$, $\|\nabla^{2}h(z)\|\leq L$,
@@
In the case of a smooth convex set, we obtain the following improvement of Corollary 3.3, which provides the convergence rate
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=O(\delta)\textrm{ as }\delta\to 0^{+}.$
@@
Let $C\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and fix $x\in\mathbb{R}^{n}\setminus C$. Assume that $\partial C$ admits a local $C^{2,1}$ chart at $\operatorname{proj}_{C}(x)$ with constants $(\rho,L,M)$. Then there exist $\mathcal{C}>0$ and $\delta_{0}>0$ such for every $0<\delta<\delta_{0}$, we have
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq$ $\,\mathcal{C}\delta+\mathcal{O}(\delta^{2}).$
@@
Consider the half-space $C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle\leq 0\}$, where $\|\nu\|=1$. Then $\partial C=\{y\in\mathbb{R}^{n}:\langle y,\nu\rangle=0\}$ is a hyperplane. Pick $x\notin C$, so that $\langle x,\nu\rangle>0$. Then
@@
Moreover, since $p_{\delta}(x)=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}\!\left[\,y\mid y\in C\,\right]$, symmetry implies that the tangential conditional mean is unchanged and the drift is purely normal. Hence,
@@
and therefore
@@ Appendix A Auxiliary lemmas
Let $D\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and for $R>0$ set $\Omega_{R}:=\operatorname{int}(D)\cap\mathbb{B}_{R}$. Then $\Omega_{R}$ is a bounded Lipschitz domain. Moreover, for a.e. $R>0$,
(8) $\mathcal{H}^{n-1}\big(\partial D\cap\partial\mathbb{B}_{R}\big)=0.$
@@
$\int_{\Omega_{R}}\operatorname{div}F(x)\,dx=\int_{\partial D\cap\mathbb{B}_{R}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial\mathbb{B}_{R}}F\cdot\nu_{\mathbb{B}_{R}}\,d\mathcal{H}^{n-1},$
where $\nu_{D}$ is the outer unit normal to $D$, defined $\mathcal{H}^{n-1}$-a.e. on $\partial D$, $\nu_{\mathbb{B}_{R}}$ is the outer unit normal to $\mathbb{B}_{R}$. Here, $F$ on $\partial\Omega_{R}$ denotes the interior Sobolev trace of $F|_{\Omega_{R}}$ on $\partial\Omega_{R}$, restricted to $\partial D\cap\mathbb{B}_{R}$ and to $\operatorname{int}(D)\cap\partial\mathbb{B}_{R}$.
@@
Moreover, $\Phi_{k}(g_{k}^{-1}(R))\subset\partial D\cap\partial\mathbb{B}_{R}$. Hence, for a.e. $R>0$,
$\mathcal{H}^{n-1}\big((\partial D\cap\partial\mathbb{B}_{R})\cap\Phi_{k}(U_{k})\big)=0.$
@@
$\partial\Omega_{R}=(\partial D\cap\mathbb{B}_{R})\cup(\operatorname{int}(D)\cap\partial\mathbb{B}_{R})\cup(\partial D\cap\partial\mathbb{B}_{R}).$
Moreover, on $\partial D\cap\mathbb{B}_{R}$ the boundary $\partial\Omega_{R}$ coincides locally with $\partial D$. Hence $\nu_{\Omega_{R}}=\nu_{D}$ $\mathcal{H}^{n-1}$-a.e. on $\partial D\cap\mathbb{B}_{R}$. Finally, by (8), the set $\partial D\cap\partial\mathbb{B}_{R}$ has $\mathcal{H}^{n-1}$-measure zero and does not contribute to the boundary integral in (9). Therefore (9) splits into
$\int_{\Omega_{R}}\operatorname{div}F\,dx=\int_{\partial D\cap\mathbb{B}_{R}}\operatorname{Tr}(F)\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial\mathbb{B}_{R}}\operatorname{Tr}(F)\cdot\nu_{\mathbb{B}_{R}}\,d\mathcal{H}^{n-1},$
@@
$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{2})\textrm{ as }\delta\to 0^{+}.$
@@
$\frac{1}{\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)]}=\frac{1}{\frac{1}{d}+\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)-\frac{1}{d}]}=d-d^{2}\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)-\frac{1}{d}]+O(\delta^{2}).$
@@
Finally, using that $|h(z)|\leq\frac{L}{2}\|z\|^{2}$ and $\mathbb{E}_{\eta_{\delta}}[\|z\|^{2}]=(n-1)\delta+O(\delta^{2})$, we obtain that
$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{2}),$
@@
$|p_{\delta}^{\nu}(x)|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{2}).$

### v3

@@ 4.1. Stochastic Approximation of Projections onto Smooth Sets
Let $C\subset\mathbb{R}^{n}$ be a closed convex set and let $p\in\partial C$. We say that $\partial C$ admits a local $C^{2,1}$ chart at $p$ with constants $(\rho,L,M)$ (see Figure 1) if, after a rigid motion sending $p$ to 0 and the outward unit normal at $p$ to $e_{n}$, there exists $\rho>0$ and a concave function $h\in C^{2,1}(\mathbb{B}^{n-1}_{\rho})$ such that $h(0)=0$, $\nabla h(0)=0$, $\|\nabla^{2}h(z)\|\leq L$,
@@
In the case of a smooth convex set, we obtain the following refinement of Corollary 3.3, which yields a better convergence rate of order $\delta$.
@@
Let $C\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and fix $x\in\mathbb{R}^{n}\setminus C$. Assume that $\partial C$ admits a local $C^{2,1}$ chart at $\operatorname{proj}_{C}(x)$ with constants $(\rho,L,M)$. Then there exists $\delta_{0}>0$ such that for every $0<\delta<\delta_{0}$, we have
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=$ $\mathcal{O}(\delta).$
Proof.
The proof can be found in Section B. ∎
Remark 4.4.
A careful review of the proof of Theorem 4.3 shows that, for this result, it suffices that the set $C$ be locally a convex body in a neighborhood of the (necessarily unique) projection point.
@@
Consider the half-space $C=\{y\in\mathbb{R}^{n}:\langle y,\nu\rangle\leq 0\}$, where $\|\nu\|=1$. Then $\partial C=\{y\in\mathbb{R}^{n}:\langle y,\nu\rangle=0\}$ is a hyperplane. Pick $x\notin C$, so that $\langle x,\nu\rangle>0$. Then
@@
Moreover, since $p_{\delta}(x)=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}\!\left[\,y\mid y\in C\,\right]$, symmetry implies that the conditional mean in tangential directions is unchanged and the drift is purely normal. Hence,
@@
where $\phi$ and $\Phi$ denote the standard normal density and distribution function, respectively. Therefore
@@
Acknowledgements
D. Morales was supported by ANID Chile under the CMM BASAL funds for the Center of Excellence FB210005 and ANID BECAS/DOCTORADO NACIONAL 21252113. P. Pérez-Aros was supported by ANID (Chile) through Fondecyt Regular grants No. 1220886, No. 1240120, and No. 1261728; CMM BASAL funds for the Center of Excellence FB210005; and the projects ECOS230027 and MATH-AmSud 23-MATH-17. E. Vilches was supported by ANID (Chile) through Fondecyt Regular grants No. 1220886, No. 1240120, and No. 1261728; CMM BASAL funds for the Center of Excellence FB210005; and the projects ECOS230027 and MATH-AmSud 23-MATH-17.
@@ Appendix A Auxiliary lemmas
Let $D\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and for $R>0$ set $\Omega_{R}:=\operatorname{int}(D)\cap B_{R}$. Then $\Omega_{R}$ is a bounded Lipschitz domain. Moreover, for a.e. $R>0$,
(6) $\mathcal{H}^{n-1}\big(\partial D\cap\partial B_{R}\big)=0.$
@@
$\int_{\Omega_{R}}\operatorname{div}F(x)\,dx=\int_{\partial D\cap B_{R}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial B_{R}}F\cdot\nu_{B_{R}}\,d\mathcal{H}^{n-1},$
where $\nu_{D}$ is the outer unit normal to $D$, defined $\mathcal{H}^{n-1}$-a.e. on $\partial D$, $\nu_{B_{R}}$ is the outer unit normal to $B_{R}$. Here, $F$ on $\partial\Omega_{R}$ denotes the interior Sobolev trace of $F|_{\Omega_{R}}$ on $\partial\Omega_{R}$, restricted to $\partial D\cap B_{R}$ and to $\operatorname{int}(D)\cap\partial B_{R}$.
@@
Moreover, $\partial D\cap\Phi_{k}(g_{k}^{-1}(R))\subset\partial D\cap\partial B_{R}$ and $\partial D\cap\partial{B}_{R^{\prime}}\cap\Phi_{k}(g_{k}^{-1}(R))=\emptyset$ for all $R^{\prime}\neq R$. Hence,
$(\partial D\cap\partial B_{R})\cap\Phi_{k}(U_{k})=(\partial D\cap\partial B_{R})\cap\Phi_{k}(g^{-1}_{k}(R)),$
which justifies that
$\mathcal{H}^{n-1}\big((\partial D\cap\partial B_{R})\cap\Phi_{k}(U_{k})\big)=0.$
@@
$\partial\Omega_{R}=(\partial D\cap B_{R})\cup(\operatorname{int}(D)\cap\partial B_{R})\cup(\partial D\cap\partial B_{R}).$
Moreover, on $\partial D\cap B_{R}$ the boundary $\partial\Omega_{R}$ coincides locally with $\partial D$. Hence $\nu_{\Omega_{R}}=\nu_{D}$ $\mathcal{H}^{n-1}$-a.e. on $\partial D\cap B_{R}$. Finally, by (6), the set $\partial D\cap\partial B_{R}$ has $\mathcal{H}^{n-1}$-measure zero and does not contribute to the boundary integral in (7). Therefore (7) splits into
$\int_{\Omega_{R}}\operatorname{div}F\,dx=\int_{\partial D\cap B_{R}}\operatorname{Tr}(F)\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial B_{R}}\operatorname{Tr}(F)\cdot\nu_{B_{R}}\,d\mathcal{H}^{n-1},$
@@
The following lemma can be proved directly using polar coordinates and integration by parts.
@@
$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{3/2})\textrm{ as }\delta\to 0^{+}.$
@@
$\frac{1}{\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)]}=\frac{1}{\frac{1}{d}+\mathbb{E}_{\eta_{\delta}}\left[g_{\delta}(z)-\frac{1}{d}\right]}=d-d^{2}\mathbb{E}_{\eta_{\delta}}\left[g_{\delta}(z)-\frac{1}{d}\right]+O(\delta^{2}).$
@@
Finally, using that $|h(z)|\leq\frac{L}{2}\|z\|^{2}$ and $\mathbb{E}_{\eta_{\delta}}[\|z\|^{2}]\leq(n-1)\delta+O(\delta^{3/2})$, we obtain that
$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{3/2}),$
@@
$|p_{\delta}^{\nu}(x)|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{3/2}).$